# Scattering of the defocusing Calogero–Moser derivative nonlinear Schrödinger equation

### v2

@@ 2.1 Generalized eigenfunctions
First we show that if $u\in H_{+}^{1}(\mathbb{R})$, the Lax operator $L_{u}$ has no eigenvalue.
@@
Let $u\in H_{+}^{1}(\mathbb{R})$, then $L_{u}$ has no eigenvalues.
@@
Let $\psi\in H_{+}^{1}(\mathbb{R})$ be an eigenvector of $L_{u}$. From the proof of [11, Proposition 5.1], we can infer
@@
Let $u\in H^{1}_{+}(\mathbb{R})$ be such that $|x|^{\alpha}u\in L^{2}(\mathbb{R})$ for some $\alpha>0$. For every $\lambda>0$, the space
@@
Let $f\in L^{\infty}$ such that $f=K_{\lambda}(f)$. Let us prove that $h\in L^{2}(\mathbb{R})$. First, we observe that
@@
If $\alpha>\frac{1}{4}$, this implies $h\in L^{2}(\mathbb{R})$. Otherwise, we plug this information in the above argument. More generally, if $|x|^{\beta}f\in L^{\infty}$ for some $\beta\geq 0$, we have $\left(1+|x|^{\alpha+\beta}\right)f\in L^{2}(\mathbb{R})$, hence, again by Hardy’s inequality,
@@
$\mathrm{e}^{i\lambda x}f(x)=\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy,$
we infer that $m\in L_{+}^{2}\text{, and finally }m\in H_{+}^{1}$. From the assumption on the kernel of $L_{u}-\lambda$, we conclude $m=0$. ∎
@@
We can also conclude the limit of $m(x,\lambda)$ as $\lambda\to 0_{+}$, which is useful in the latter proof.
@@
Let $\ell_{m}(-\infty)=1$. For any $x\in\mathbb{R}$, we have
@@
$\left|m(x,\lambda)-1\right|\leq\left|(\operatorname{Id}-K_{\lambda})^{-1}1-(\operatorname{Id}-K_{0})^{-1}1)\right|\rightarrow 0\text{ as }\lambda\to 0_{+}.$
@@ 2.2 The distorted Fourier transform
From Proposition 2.4, for $u\in H_{+}^{1}(\mathbb{R}),|x|^{\alpha}u\in L^{2}(\mathbb{R})$ with some $\alpha>0$, there exists a unique $m_{-}\in\mathscr{M}_{u}(\lambda)$ such that $\ell_{m_{-}}(-\infty)=1$. Given $f\in L_{+}^{1}(\mathbb{R})$, we can define
$\mathcal{F}_{d}f(\lambda):=\widetilde{f}(\lambda):=\int_{\mathbb{R}}f(x)\overline{m_{-}(x,\lambda)}dx,\lambda\geq 0.$
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}([0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}(L_{u})$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
@@ 2.3 Computation of the scattering term
We keep the assumption $u\in\mathcal{S}_{+}(\mathbb{R})$. Let $\widetilde{f}\in H^{1}(0,\infty)$, $f\in\operatorname{Dom}(X^{*})$ and $f\in H_{+}^{1}(\mathbb{R})$, then
@@
Here $\chi_{\varepsilon}(\lambda)=\mathrm{e}^{-\varepsilon\lambda^{2}}$, and the second identity comes from the integral by parts and Lemma 2.7.
@@
Then from the definition of $X^{*}$ in (1.5) and the identity (1.7), we can deduce that $\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf=X^{*}f$ and
@@
$\widetilde{h}(\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta.$
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty),h\in H_{+}^{1}(\mathbb{R})$ and $h\in\operatorname{Dom}(X^{*})$ using the Schur test, so we can apply (2.11) for $h$. Now we justify $I_{+}(h)$. From (2.10), we have
@@
where $\widehat{v}(\xi)=\widetilde{u}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$. From the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$ and (2.11), we can obtain (2.12) for $u\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$.
@@
In fact, we can easily see that the scattering term in Theorem 1.5 is exacly $\mathrm{e}^{it\partial_{x}^{2}}v$ obtained in (2.12) with $u=u_{0}$.
@@ 3 Long-time behavior
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u$ to the defocusing (CM-DNLS) (1.1) satisfies
@@
Since $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})$, we infer $u_{0}\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$. From Corollary 2.13, we know that $\widetilde{u}_{0}\in L^{4}(0,\infty)$ and
$\|\widetilde{u}_{0}\|_{L^{4}(0,\infty)}\leq C\|u\|_{L^{\frac{4}{3}}(\mathbb{R})}.$ (3.2)
The mass conservation law $\|u(t)\|_{L^{2}(\mathbb{R})}=\|u_{0}\|_{L^{2}(\mathbb{R})}$ implies that u(t){u(t)} is uniformly bounded in $L_{+}^{2}(\mathbb{R})$, hence it admits weakly convergent subsequences in $L_{+}^{2}(\mathbb{R})$ as $t\to\infty$. We shall prove that the weak limit is unique.
Let us recall the explicit formula of solution to the defocusing (CM-DNLS) (1.1),
$\forall z\in\mathbb{C}_{+},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (3.3)
Now we denote
$\mathcal{F}_{d}^{-1}(i\partial_{\lambda}\widetilde{f}):=Yf,\hskip 18.49988pt\forall\widetilde{f}\in H^{1}(0,\infty).$
We recall (2.12) and we rewrite it as
$h_{0}:=(Y+2tL_{u_{0}}-zId)^{-1}u_{0}$
and
$\frac{1}{2i\pi}I_{+}\left[(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right]=\frac{1}{2i\pi}I_{+}\left[h_{0}\right]=\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z).$
Then we have
$\,u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)$
$=$ $\,\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right]-\frac{1}{2i\pi}I_{+}\left[(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right]$
$=$ $\,\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}(Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right].$
The last identity above comes from the second resolvent identity.
Then from Lemma 1.4, (2.11) and (3.2), we infer
$|u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)|$
$\leq\,$ $C(z)\left\|(Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right\|_{L^{2}(\mathbb{R})}$
$\leq\,$ $C(z)\left\|\mathcal{F}_{d}\left((Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right)\right\|_{L^{2}(0,\infty)}$
$\leq\,$ $C(z)\left\|i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{h}_{0})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{h}_{0}(\lambda)\right\|_{L^{2}(0,\infty)}$
$\leq\,$ $C(z)\|\widetilde{u_{0}}\|_{L^{4}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)}$
$\leq\,$ $C(z)\|u_{0}\|_{L^{\frac{4}{3}}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)},$
where
$\widetilde{h_{0}}(t,\lambda)=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}_{0}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u_{0}}(\eta)d\eta.$
@@
Following the approach developed in Section 2.3 and Section 3, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$.
We fix $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. From [11, Section 5], we know that every eigenvalue of $L_{u_{0}}=D-u_{0}T_{\bar{u}_{0}}$ (the Lax operator of (1.3)) is simple, and the number $N$ of eigenvalues of $L_{u_{0}}$ is finite with $N\leq\frac{\|u_{0}\|_{L^{2}}^{2}}{2\pi}$. From [10, Theorem 4.3], we know that for every $f\in L_{+}^{2}(\mathbb{R})$
@@
Here $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, $\mathcal{F}_{d}(f)=\widetilde{f}$ is the distorted Fourier transform defined in [10, (4.1)], $P_{p}(L_{u_{0}})$ denote the projection onto the eigensubspace of $L_{u_{0}}$ and $P_{ac}(L_{u_{0}})$ denote the projection onto the absolutely continuous subspace of $L_{u_{0}}$. Now we apply (3.4) for $f=u_{0}$, and we obtain
@@
Now we assume that the solution $u(t,\cdot)\in H_{+}^{1}(\mathbb{R})$ to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}$ exists globally in time (this is always true if $\|u_{0}\|_{L^{2}}^{2}\leq 2\pi$). From the explicit formula (3.3), we infer that for every $t\in\mathbb{R}$ and every $z\in\mathbb{C}_{+}$,
@@
where $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$. From this we can deduce the sacttering term of the focusing (CM-DNLS) (1.3). Notice that if $\|u_{0}\|_{L^{2}}^{2}<2\pi$, $L_{u_{0}}$ has no eigenvalues. In this case $P_{p}(L_{u_{0}})=0$ and $P_{ac}(L_{u_{0}})u_{0}=u_{0}$, so from the above argument we can directly deduce the scattering of the focusing (CM-DNLS) (1.3) for $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with $\alpha>\frac{1}{4}$.
Now we analyse the first term on the right hand side in (3.6). Following the method used in the proof of [8, Lemma 3.1], we can also deduce that

### v3

@@ 2.1 Generalized eigenfunctions
First we show that if $u\in L_{+}^{2}(\mathbb{R})$, the Lax operator $L_{u}$ has no eigenvalue.
@@
Let $u\in L_{+}^{2}(\mathbb{R})$, then $L_{u}$ has no eigenvalues.
@@
Let $\psi\in H_{+}^{\frac{1}{2}}(\mathbb{R})$ be an eigenvector of $L_{u}$. From the proof of [12, Proposition 5.1], we can infer
@@
Let $u\in L^{2,\alpha}_{+}(\mathbb{R})$ with some $\alpha>0$. For every $\lambda>0$, the space
@@
Let $f\in L^{\infty}$ such that $f=K_{\lambda}(f)$. Let us prove that $f\in L^{2}(\mathbb{R})$. First, we observe that
@@
If $\alpha>\frac{1}{4}$, this implies $f\in L^{2}(\mathbb{R})$. Otherwise, we plug this information in the above argument. More generally, if $|x|^{\beta}f\in L^{\infty}$ for some $\beta\geq 0$, we have $\left(1+|x|^{\alpha+\beta}\right)f\in L^{2}(\mathbb{R})$, hence, again by Hardy’s inequality,
@@
$m(x)=\mathrm{e}^{i\lambda x}f(x),$
we infer that $m\in L_{+}^{2}\text{, and finally }m\in H_{+}^{\frac{1}{2}}$ is an eigenvector of $L_{u_{0}}$. Since the kernel of $L_{u}-\lambda$ is trivial, we conclude $m=0$. ∎
@@
If $u\in\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$, we can also conclude the limit of $m(x,\lambda)$ as $\lambda\to 0_{+}$, which is useful in the latter proof.
@@
Let $u\in\mathcal{S}_{+}(\mathbb{R})$ and $\ell_{m}(-\infty)=1$. For any $x\in\mathbb{R}$, we have
@@
$\left\|a(\cdot,\lambda)-1\right\|_{L^{\infty}}\leq\left\|(\operatorname{Id}-K_{\lambda})^{-1}1-(\operatorname{Id}-K_{0})^{-1}1)\right\|_{L^{\infty}}\rightarrow 0\text{ as }\lambda\to 0_{+}.$
Since for any $x\in\mathbb{R}$, $\mathrm{e}^{i\lambda x}\to 1$ as $\lambda\to 0_{+}$, we can conclude the lemma. ∎
@@ 2.2 The distorted Fourier transform
From Proposition 2.4, for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, there exists a unique $m_{-}\in\mathscr{M}_{u}(\lambda)$ such that $\ell_{m_{-}}(-\infty)=1$. Given $f\in L_{+}^{1}(\mathbb{R})$, we can define
$\mathcal{F}_{d}f(\lambda):=\widetilde{f}(\lambda):=\int_{\mathbb{R}}f(x)\overline{m_{-}(x,\lambda)}dx,\qquad\lambda>0.$
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}((0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}(L_{u})$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
@@
Remark 2.10.
By the polarization identity, the isometry property implies preservation of the inner product, i.e., for any $f,g\in L_{+}^{2}(\mathbb{R})$, we have
$\left\langle f,g\right\rangle_{L^{2}(\mathbb{R})}=\frac{1}{2\pi}\langle\widetilde{f},\widetilde{g}\rangle_{L^{2}(0,\infty)}.$ (2.4)
@@ 2.3 Computation of the scattering term
We keep the assumption $u\in\mathcal{S}_{+}(\mathbb{R})$. Let $\widetilde{f}\in H^{1}(0,\infty)$ and $f\in H_{+}^{1}(\mathbb{R})$, then
@@
Here $\chi_{\varepsilon}(\lambda)=\mathrm{e}^{-\varepsilon\lambda^{2}}$, and the second identity comes from the integration by parts and Lemma 2.7.
@@
Then from the definition of $X^{*}$ in (1.5) and the identity (1.7), we can deduce that $f\in\operatorname{Dom}(X^{*})$, $\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf=X^{*}f$ and
@@
$\widetilde{h}(t,\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta.$ (2.13)
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty)$ and $h\in H_{+}^{1}(\mathbb{R})$ using the Schur test, so we can apply (2.12) for $h$. Now we justify $I_{+}(h)$. From (2.11), we have
@@
where $\widehat{v}(\xi)=\widetilde{u}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$.
@@
In fact, we can easily see that the scattering term in Theorem 1.9 is exacly $\mathrm{e}^{it\partial_{x}^{2}}v$ obtained in (2.14) with $u=u_{0}$. We remark that, from the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{2}(\mathbb{R})$, we can obtain (2.13) for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
2.4 Extension of the formula
Let $u\in\mathcal{S}_{+}(\mathbb{R})$. From the previous section, we know that for $\widetilde{f}\in H^{1}(0,\infty)$ and $f\in H_{+}^{1}(\mathbb{R})$, we have $f\in\operatorname{Dom}(X^{*})$ with
$I_{+}(f)=J_{+}(\widetilde{f}):=\widetilde{f}(0_{+})$ (2.15)
and
$\mathcal{F}_{d}\left(X^{*}f\right)=i\partial_{\lambda}\widetilde{f}-\frac{1}{4\pi}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right):=i\partial_{\lambda}\widetilde{f}+B(\widetilde{u})\widetilde{f}$. (2.16)
Furthermore, we know that, for every $\widetilde{f}\in L^{2}(0,\infty)$,
$(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{f}=\mathrm{e}^{it\lambda^{2}}(i\partial_{\lambda}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}\widetilde{f}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{f}(\eta)d\eta.$
Using the Schur test, we can easily see that $(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{f}\in L^{2}(0,\infty)$, so $(i\partial_{\lambda}+2t\lambda-z)^{-1}$ is well-defined on $L^{2}(0,\infty)$. It is also not hard to see that $\partial_{\lambda}$ is a dissipative operator, therefore $\partial_{\lambda}$ is a maximally dissipative operator. Since B(u~)B($\widetilde${u}) is a self-adjoint operator, iB(u~)-iB($\widetilde${u}) is a dissipative operator. In fact, we can easily observe that
$(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}=\mathrm{e}^{it\lambda^{2}}(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}.$ (2.17)
As $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}$ is dissipative and bounded on $L^{2}(0,\infty)$, from the Miyadera–Gustafson theorem 1.5, we can infer that $(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}$ is well-defined on $L^{2}(0,\infty)$. It is not hard to see that
$\widetilde{g}:=(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}\widetilde{u}$
satisfying $\widetilde{g}\in H^{1}(0,\infty)$ and $g\in H_{+}^{1}(\mathbb{R})$, then from (2.15) and (2.16), we have
$J_{+}(\widetilde{g})=I_{+}(g),$
where
$g=\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u.$
Then from Theorem 1.8, we can deduce that, if $u(t)$ is a solution of (1.1) with $u(0)=u_{0}\in\mathcal{S}_{+}(\mathbb{R})$, then for every $z\in\mathbb{C}_{+}$,
$u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right]=\frac{1}{2i\pi}J_{+}\left[(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z)^{-1}\widetilde{u}_{0}\right].$ (2.18)
Now we compute $u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)$ with $\widehat{v}_{0}(\xi)=\widetilde{u}_{0}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$. From (2.14), (2.18) and the second resolvent identity, we can infer that
$u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)=\frac{1}{2i\pi}J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right].$ (2.19)
In fact, it is not hard to see that $\partial_{\lambda}$ is also a maximally disspative operator on $L^{1}(0,\infty)$ with $D(\partial_{\lambda})=W^{1,1}(0,\infty)$. For $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, we can infer that $u_{0}\in L_{+}^{p}(\mathbb{R})$ with some $1<p<2$. Then from Corollary 2.14, we can deduce that $\widetilde{u_{0}}\in L^{q}$ with some $2<q<\infty$. Then for $f\in W^{1,1}(0,\infty)$, we have
$\left\|\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}f\right\|_{L^{1}(0,\infty)}\leq C\|\widetilde{u}_{0}\|_{L^{q}}^{2}\|f\|_{L^{r}}\leq C\|\widetilde{u}_{0}\|_{L^{q}}^{2}\|f\|_{L^{1}}^{\frac{1}{r}}\|f^{\prime}\|_{L^{1}}^{1-\frac{1}{r}},$
here $r$ satisfies $1<r<\infty$. It is not hard to see that $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}$ is a dissipative operator on $L^{1}(0,\infty)$. Combine the above estimate with Young’s inequality for products, we can deduce that $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}$ is infinitesimally small with respect to $\partial_{\lambda}$. Then from the Miyadera–Gustafson theorem 1.5, we can infer that $\partial_{\lambda}-i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}$ is a maximally dissipative operator and $(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}$ is well-defined on $L^{1}(0,\infty)$. More precisely, we can deduce that, for any $f\in L^{1}(0,\infty)$,
$(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}f\in W^{1,1}(0,\infty)$
with
$\left\|(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}f\right\|_{W^{1,1}(0,\infty)}\leq C\left\|f\right\|_{L^{1}(0,\infty)}.$
Therefore we have
$\left|J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$ (2.20)
$=$ $\,\left|J_{+}\left[(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$
$\leq$ $\,C\left\|(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{W^{1,1}(0,\infty)}$
$\leq$ $\,C\left\|B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{1}(0,\infty)}$
$\leq$ $\,C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\left\|\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{\infty}(0,\infty)}$
$\leq$ $\,C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{3}$
$\leq$ $\,C\left\|u_{0}\right\|_{L^{2}(\mathbb{R})}.$
Then by the density from $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{2}(\mathbb{R})$, we can obtain (2.19) and (2.20) for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
@@ 3 Long-time behavior
Let $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R}):=\{w\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. Then the solution $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
@@
First of all, the mass conservation law $\|u(t)\|_{L^{2}(\mathbb{R})}=\|u_{0}\|_{L^{2}(\mathbb{R})}$ implies that u(t){u(t)} is uniformly bounded in $L_{+}^{2}(\mathbb{R})$, hence it admits weakly convergent subsequences in $L_{+}^{2}(\mathbb{R})$ as $t\to\infty$. We shall prove that the weak limit is unique.
Let us recall the formula (2.19) for $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$,
$u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)=\frac{1}{2i\pi}J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right].$ (3.2)
We recall (2.13) for $u=u_{0}$ and we rewrite it as
$\widetilde{h}_{0}(t,\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}_{0}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}_{0}(\eta)d\eta.$
Then from (2.20) and (3.2), we infer
$\left|u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)\right|$
$\leq\,$ $C\left|J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$
$\leq\,$ $C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\left\|\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{\infty}(0,\infty)}$
$\leq\,$ $C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L^{\infty}(0,\infty)}.$
@@
Following the approach developed in Section 2.3 and Section 3, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for some $\alpha>0$.
We fix $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ for some $\alpha>0$. From [12, Section 5], we know that every eigenvalue of $L_{u_{0}}=D-u_{0}T_{\bar{u}_{0}}$ (the Lax operator of (1.3)) is simple, and the number $N$ of eigenvalues of $L_{u_{0}}$ is finite with $N\leq\frac{\|u_{0}\|_{L^{2}}^{2}}{2\pi}$. From [11, Theorem 4.3], we know that for every $f\in L_{+}^{2}(\mathbb{R})$
@@
Here $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, $\mathcal{F}_{d}(f)=\widetilde{f}$ is the distorted Fourier transform defined in [11, (4.1)], $P_{p}(L_{u_{0}})$ denote the projection onto the eigensubspace of $L_{u_{0}}$ and $P_{ac}(L_{u_{0}})$ denote the projection onto the absolutely continuous subspace of $L_{u_{0}}$. Now we apply (3.3) for $f=u_{0}$, and we obtain
@@
Now we assume that the solution $u(t,\cdot)\in L_{+}^{2}(\mathbb{R})$ to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}$ exists globally in time (this is always true if $\|u_{0}\|_{L^{2}}^{2}<2\pi$ or if $u_{0}\in H_{+}^{1}(\mathbb{R})$ with $\|u_{0}\|_{L^{2}}^{2}=2\pi$). From the explicit formula (3.2), we infer that for every $t\in\mathbb{R}$ and every $z\in\mathbb{C}_{+}$,
@@
where $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$. From this we can deduce the sacttering term of the focusing (CM-DNLS) (1.3). Notice that if $\|u_{0}\|_{L^{2}}^{2}<2\pi$, $L_{u_{0}}$ has no eigenvalues. In this case $P_{p}(L_{u_{0}})=0$ and $P_{ac}(L_{u_{0}})u_{0}=u_{0}$, so from the above argument we can directly deduce the scattering of the focusing (CM-DNLS) (1.3) for $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with some $\alpha>0$.
Now we analyse the first term on the right hand side in (3.5). Following the method used in the proof of [9, Lemma 3.1], we can also deduce that